OpenStudy (anonymous):

given f(x)=(x^5)+3 and g(x)=5rootx-2 find (f*g)(x)

OpenStudy (anonymous):

f(x)=\[x^{5}\]+3 and g(x)=\[\sqrt[5]{x-2?}\]

OpenStudy (radar):

Assuming this is asking for the composite f(g(x)).

OpenStudy (anonymous):

yes

OpenStudy (radar):

f(g(x))=[g(x)]^5=[(x-2^1/5)]^5

OpenStudy (radar):

\[[(x-2)^{1/5}]^{5}\]=x-2

OpenStudy (radar):

Does that look right ???? Someone elses input??

OpenStudy (anonymous):

lost me on the 1/5

OpenStudy (anonymous):

Looks right to me but you dropped the +3 :)

OpenStudy (radar):

\[x ^{1/5}=\sqrt[5]{x}\]

OpenStudy (anonymous):

wait never mind got you thanks alot

OpenStudy (anonymous):

So your final answer should be x+1.

OpenStudy (radar):

ooops didn't see the +3,that makes a difference. lol

OpenStudy (anonymous):

Not by much. You showed the hard part :)

OpenStudy (anonymous):

thank you both

OpenStudy (radar):

Thanks, just reviewed this about a week ago (composites) still learning!

OpenStudy (anonymous):

My brain is rusty on most of this. Don't feel bad. It always feels by the end of a class that the stuff from the beginning has been pushed out ;)

OpenStudy (radar):

Yes, FIFO, lol